OpenStudy (anonymous):

write an equation of a line that satifies the following conditions: passes through (-1,-7) and (1,3)

OpenStudy (fanduekisses):

Use the point slope formula

OpenStudy (fanduekisses):

and the slope form

OpenStudy (fanduekisses):

use slope formula to find the slope \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (fanduekisses):

soslope: \[\frac{ 3- -7 }{ 1- -1 }=5\]

OpenStudy (fanduekisses):

Now point slope formula: y – y1 = m(x – x1), so y--7=5(x-1) = y=5x-12

OpenStudy (fanduekisses):

\[y=5x-12\]

OpenStudy (fanduekisses):

is the equation of the line :)